# Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations–B Atmospheric Characterisation with the Twinkle Space Telescope Following Advances from JWST Observations

### v1

@@ 3.1 Model assumptions: lessons learned from JWST
| Planet | Median SNR | Predicted Transits (5) | Predicted Transits (3) | Predicted Transits (2) | Realistic Transits |
| HD 209458 b | 5 | 1 (0.18) | 1 (0.51) | 2 (1.21) | 1 (0.85) |
| 7 | 1 (0.35) | 2 (1.04) | 3 (2.6) | 2 (1.77) | |
| 10 | 1 (0.75) | 3 (2.32) | 7 (6.79) | 5 (4.24) | |
| WASP-39 b | 5 | 3 (2.26) | 7 (6.31) | 15 (14.31) | 3 (2.11) |
| 7 | 5 (4.45) | 13 (12.45) | 29 (28.45) | 5 (4.14) | |
| 10 | 10 (9.12) | 26 (25.74) | 60 (59.84) | 9 (8.48) | |
| WASP-80 b | 5 | 3 (2.01) | 6 (5.7) | 14 (13.41) | 8 (7.49) |
| 7 | 4 (3.98) | 12 (11.56) | 29 (28.5) | 16 (15.35) | |
| 10 | 9 (8.33) | 26 (25.44) | 71 (70.86) | 35 (34.67) | |
| WASP-107 b | 5 | 1 (0.16) | 1 (0.45) | 2 (1.02) | 2 (1.31) |
| 7 | 1 (0.32) | 1 (0.88) | 2 (2.00) | 3 (2.59) | |
| 10 | 1 (0.65) | 2 (1.81) | 5 (4.15) | 6 (5.4) | |
| GJ 3470 b | 3 | 1 (0.59) | 2 (1.65) | 4 (3.77) | 4 (3.58) |
| 5 | 2 (1.64) | 5 (4.69) | 12 (11.09) | 11 (10.49) | |
| 7 | 4 (3.27) | 10 (9.55) | 24 (23.85) | 23 (22.45) |
@@ 3.2 Candidate List for Atmospheric Characterisation Survey
This initial candidate list used for the observability study for Twinkle’s atmospheric characterisation survey is specifically tailored to H2/He-rich planets, due to current uncertainties surrounding the atmospheric composition of smaller planets. Recent observations of terrestrial planets have highlighted significant diversity, emphasising the need for further study with dedicated future missions. Consequently, terrestrial planets have not been extensively analysed in this observability study. Further refinement of the candidate list will be informed by complementary observations from JWST, PLATO, CSST, and Earth 2.0, enabling tailored observational strategies aligned with the scientific objectives of the Twinkle Science Team.
@@ 3.3.1 HD 209458 b
To provide a more comprehensive assessment, we performed a free retrieval analysis without assuming chemical equilibrium by selecting the eleven most critical molecules identified in previous simulations, along with species suggested by Xue et al. (2024). A constant vertical mixing ratio (VMR) profile based on equilibrium abundances and suggestions from Xue et al. (2024) was used for these molecules in a free retrieval analysis. Results in Figure 11 confirm robust retrieval of H2O, TiO, and VO, whereas other molecular signatures remain masked by the dominant H2O absorption and cloud opacity. Twinkle’s extensive spectral coverage from optical to infrared allows for clear identification of TiO and VO absorption features. However, due to the relatively low assumed abundance of CO2 () and increased noise at longer wavelengths, the CO2 feature was only partially detected and thus less reliably constrained.
@@
| Parameter | Truth | Retrieved (SNR = 5) |
| Planet Radius [] | 1.35 | |
| Temperature [K] | 1140 | |
| Detection | ||
| Upper Limit | ||
| 2.11 | ||
@@ 3.3.2 WASP-107 b
| Parameter | Truth | RP (SNR = 5) |
| Planet Radius [] | 0.933 | |
| [K] | 700 | |
| Detection | ||
| Upper Limit | ||
@@ 3.3.3 GJ 3470 b
| Parameter | Truth | RP (SNR = 3) |
| Planet Radius [] | 0.36 | |
| Temperature [K] | 600 | |
| Detection | ||
| Upper Limit | ||
@@ 3.4 Spectral Retrievals: super-Earths
Figure 7 shows a simulated spectrum for 55 Cnc e, along with Twinkle’s predicted error after fitting with available JWST NIRCam data from Hu et al. (2024). This baseline scenario uses three blackbody curves representing temperatures of 1400, 2000, and 2500 K. Comparing these baselines with the simulated and observed data points reveals distinctive absorption features present in the emission spectrum. Twinkle’s expected performance, calculated by reducing the resolution of channel 0 to one-third and channel 1 to half their native resolutions and stacking ten eclipse observations, demonstrates comparable sensitivity to JWST within its spectral range. A combined analysis of Figures 7 and 8 reveals significant spectral divergence from blackbody emission beginning around 2.0 m. Although the higher resolution of channel 0 below this wavelength is underutilised, the spectral range above 2.0 m effectively captures prominent CO2 absorption features. Additional CO2 features are clearly identifiable within Twinkle’s channel 1 coverage. Figure 8 further highlights that CO2 dominates the spectral signature throughout Twinkle’s wavelength range, overshadowing contributions from minor species such as C2H2, CO, and HCN. Notably, an increase in CO contribution is evident beyond 4.2 m, suggesting enhanced detectability if Twinkle’s coverage could extend slightly further.
Using this simulated spectrum as input, we performed emission retrievals to evaluate Twinkle’s capability in constraining atmospheric properties. To prevent spectral bias from wavelength binning, retrievals were conducted at Twinkle’s native resolution, despite larger resultant uncertainties. Retrieval outcomes for scenarios involving 10, 20, and 30 stacked eclipse observations are presented in Figure 14. The results illustrate that, with just 10 eclipses, Twinkle effectively constrains the temperature structure and reliably retrieves the abundance of the dominant atmospheric species, CO2. Additional observations enhance parameter constraints incrementally but do not significantly improve the detection of dominant atmospheric constituents. This demonstrates Twinkle’s robust performance, particularly for bright terrestrial targets.
| Parameter | Truth | RP (10 eclipses) |
| Planet Radius [] | 0.167 | |
| [K] | 2300 | |
| [K] | 1500 | |
| [Pa]) | 5 | |
| [Pa]) | 2 | |
| Detection | ||
| Upper Limit | ||
@@ 3.5 Minor Species Sensitivity Study on WASP-107 b
Contribution plots from simulations (Figures 5, 6, and 8) revealed that several minor species, such as CH4, SO2, NH3, and H2S, typically possess abundances too low to yield distinguishable spectral features, especially in atmospheres dominated by water vapour absorption or obscured by clouds. Direct retrieval of these species from featureless continuum regions was thus found unreliable, as illustrated in posterior plots provided in the Appendix. Such limitations reflect realistic observational challenges faced when attempting to detect minor constituents even at higher SNR levels.
To better quantify these observational constraints, we assessed the amplification factors necessary for each minor species’ spectral features to surpass the noise threshold of Twinkle observations at SNR = 10 (Comparison of the original and amplified sepctra in Figure 9). Our results show that CH4 and SO2 features become clearly observable when their abundances increase by approximately fivefold, corresponding to abundances of and , respectively. For NH3 and H2S, which have broader and weaker absorption bands overlapping significantly with water vapour, spectral signatures required approximately a twentyfold abundance increase (to and ) before becoming discernible.
@@ 4.1 H2/He dominated exoplanets
In summary, Twinkle’s capability to characterise exoplanet atmospheres depends primarily on stellar brightness in optical and infrared wavelengths, planetary and stellar radii, main atmospheric component and cloud coverage. While stellar brightness and planetary dimensions are usually well-characterised by previous observations, cloud properties and, in the case of small planets, atmospheric mean molecular weight may require dedicated characterisation efforts. Cloud contributions may significantly impact parameters’ retrieval, especially for minor species. For instance, in WASP-107 b, SO2 spectral features were fully obscured by cloud opacity, and there are degeneracies between cloud height and planetary radius. Incorporating detailed cloud microphysical and radiative models into future simulations might allow more realistic analyses.
@@ 4.3 Stellar Activity Effects, Targets Observability and Future Works
While stacking multiple observations to achieve higher SNR per wavelength bin is an effective strategy to enhance target observability, variability between observations introduces potential challenges. Variations in planetary atmospheric conditions between different transits or eclipses can lead to differences in the collected spectra. More significantly, stellar activity may significantly contaminate atmospheric parameters such as molecular abundances and temperature retrievals. A population study of 20 exoplanets conducted with HST by Saba et al. (2025) demonstrated that stellar contamination could alter molecular abundance estimates by up to six orders of magnitude and atmospheric temperatures by up to 145%. Stellar contamination predominantly affects the optical wavelengths, which are partially covered by Twinkle. Hence, the potential influence of stellar activity must be carefully considered, especially when multiple observations are stacked. Investigating stellar activity effects on exoplanet atmospheres is one of Twinkle’s core scientific themes, and a future study will explore Twinkle’s capability to identify and quantify these stellar variations.
@@ 5 Conclusions
Our analysis indicates that approximately 100 exoplanets are suitable for in depth atmospheric studies and over 300 exoplanets are good candidates for characterisation.
@@ Acknowledgements
We gratefully acknowledge insightful discussions and inputs from members of the Twinkle Science Team, and their participation in Twinkle Science Conference.
@@
Appendix A Retrieval Posterior Plots
@@
Appendix B Mass Estimates for TESS Object of Interests
In order to estimate masses for TOIs with radius available, we utilised Forecaster tool developed by Chen & Kipping (2017), a code that predict mass from radius for objects. We take the maximum radius of known exoplanets as an edge, and treat those with larger radii as dwarfs and ignore them in this study. Mass estimate using Forecaster tool for TOIs with radius less than 8 Earth radii is being adopted as they align with the known exoplanets distribution as shown in Figure 15. For TOIs with higher radius, we assume an asymmetric log-normal distribution to randomly set the mass, with the mean value set as the median value of known exoplanets with radius larger than 8 Earth radii. Resultant mass distribution for TOIs in Twinkle’s FOR is shown in Figure 15.

### v2

@@ 3.1 Model assumptions: lessons learned from JWST
[…]
@@ 3.2 Candidate List for Atmospheric Characterisation Survey
This initial candidate list used for the observability study for Twinkle’s atmospheric characterisation survey is specifically tailored to H2/He-rich planets, due to current uncertainties surrounding the atmospheric composition of smaller planets. Recent observations of terrestrial planets have highlighted significant diversity, emphasising the need for further study with dedicated future missions. Consequently, terrestrial planets have not been extensively analysed in this observability study. Further refinement of the candidate list will be informed by complementary observations from JWST, PLATO, CSST, and Earth 2.0, enabling tailored observational strategies aligned with the scientific objectives of the Twinkle Science Team. The candidate list presented in this work will be made publicly available to support the broader exoplanet community in planning complementary observations.
@@ 3.3.1 HD 209458 b
To provide a more comprehensive assessment, we performed a free retrieval analysis without assuming chemical equilibrium by selecting the eleven most critical molecules identified in previous simulations, along with species suggested by Xue et al. (2024). A constant vertical mixing ratio (VMR) profile based on equilibrium abundances and suggestions from Xue et al. (2024) was used for these molecules in a free retrieval analysis. Results in Figure 12 confirm robust retrieval of H2O and TiO, whereas other molecular signatures remain masked by the dominant H2O absorption and cloud opacity. Twinkle’s extensive spectral coverage from optical to infrared allows for clear identification of TiO absorption features, while VO remains below the detection threshold (LRT ; Table 4) despite showing a posterior shift from the prior. However, due to the relatively low assumed abundance of CO2 () and increased noise at longer wavelengths, the CO2 feature was only partially detected and thus less reliably constrained.
@@
| Parameter | Truth | Retrieved (SNR = 5) | LRT () |
| Planet Radius [] | 1.35 | ||
| Temperature [K] | 1140 | ||
| Detection | |||
| 15.18 | |||
| 8.99 | |||
| Upper Limit | |||
| 2.05 | |||
| 1.20 | |||
| 0.10 | |||
| 0.05 | |||
| 0.12 | |||
| 2.11 | |||
@@ 3.3.2 WASP-107 b
| Parameter | Truth | RP (SNR = 5) | LRT () |
| Planet Radius [] | 0.933 | ||
| Temperature [K] | 700 | ||
| Detection | |||
| 13.39 | |||
| Upper Limit | |||
| 1.35 | |||
| 0.07 | |||
| 0.08 | |||
| 0.01 | |||
@@ 3.3.3 GJ 3470 b
| Parameter | Truth | RP (SNR = 3) | LRT () |
| Planet Radius [] | 0.36 | ||
| Temperature [K] | 600 | ||
| Detection | |||
| 5.03 | |||
| 3.37 | |||
| 9.52 | |||
| Upper Limit | |||
| 0.02 | |||
| 0.39 | |||
@@ 3.4 Spectral Retrievals: super-Earths
Figure 7 shows a simulated spectrum for 55 Cnc e, along with Twinkle’s predicted error after fitting with available JWST NIRCam data from Hu et al. (2024). This baseline scenario uses three blackbody curves representing temperatures of 1400, 2000, and 2500 K. Comparing these baselines with the simulated and observed data points reveals distinctive absorption features present in the emission spectrum. Twinkle’s expected performance, calculated by reducing the resolution of channel 0 to one-third and channel 1 to half their native resolutions and stacking ten eclipse observations, demonstrates comparable sensitivity to JWST within its spectral range. Figure 7 reveals significant spectral divergence from blackbody emission beginning around 2.0 m. Although the higher resolution of channel 0 below this wavelength is underutilised, the spectral range above 2.0 m effectively captures prominent CO2 absorption features. Additional CO2 features are clearly identifiable within Twinkle’s channel 1 coverage. The contribution curves further highlight that CO2 dominates the spectral signature throughout Twinkle’s wavelength range, overshadowing contributions from minor species such as C2H2, CO, and HCN. Notably, an increase in CO contribution is evident beyond 4.2 m, suggesting enhanced detectability if Twinkle’s coverage could extend slightly further.
Using this simulated spectrum as input, we performed emission retrievals to evaluate Twinkle’s capability in constraining atmospheric properties. To avoid the loss of spectral information introduced by post-processing wavelength binning, retrievals were conducted at Twinkle’s native resolution, despite larger resultant uncertainties. Retrieval outcomes for scenarios involving 10, 20, and 30 stacked eclipse observations are presented in Figure 15. The results illustrate that, with just 10 eclipses, Twinkle effectively constrains the temperature structure and reliably retrieves the abundance of the dominant atmospheric species, CO2. Additional observations enhance parameter constraints incrementally but do not significantly improve the detection of dominant atmospheric constituents. This demonstrates Twinkle’s robust performance, particularly for bright super-Earth targets such as 55 Cnc e.
| Parameter | Truth | RP (10 eclipses) | LRT () |
| Planet Radius [] | 0.167 | ||
| [K] | 2300 | ||
| [K] | 1500 | ||
| [Pa]) | 5 | ||
| [Pa]) | 2 | ||
| Detection | |||
| 5.38 | |||
| Upper Limit | |||
| 0.01 | |||
| 0.01 | |||
@@ 3.5 Minor Species Sensitivity Study on WASP-107 b
Contribution plots from simulations (Figures 5, 6, and 7) revealed that several minor species, such as CH4, SO2, NH3, and H2S, typically possess abundances too low to yield distinguishable spectral features, especially in atmospheres dominated by water vapour absorption or obscured by clouds. Direct retrieval of these species from featureless continuum regions was thus found unreliable, as illustrated in posterior plots provided in the Appendix. Such limitations reflect realistic observational challenges faced when attempting to detect minor constituents even at higher SNR levels.
To better quantify these observational constraints, we assessed the amplification factors necessary for each minor species’ spectral features to surpass the noise threshold of Twinkle observations at SNR = 10 (Comparison of the original and amplified spectra in Figure 8). Our results show that CH4 and SO2 features become clearly observable when their abundances increase by approximately fivefold, corresponding to abundances of and , respectively. For NH3 and H2S, which have broader and weaker absorption bands overlapping significantly with water vapour, spectral signatures required approximately a twentyfold abundance increase (to and ) before becoming discernible.
@@
While the amplified spectra may appear visually consistent with the baseline within individual spectral bin error bars (Figure 8), the LRT analysis captures the cumulative effect of systematic spectral differences across multiple bins simultaneously. The retrieval posteriors in Figure 9 confirm that the amplified species are indeed constrained, validating the sensitivity thresholds identified by our analysis.
| Molecule | Amplification | SNR 10 | SNR 5 |
| SO2 | baseline | ||
| 5 | 4.69 | 0.12 | |
| 9 | – | 6.20 | |
| NH3 | baseline | 0.01 | – |
| 5 | 0.03 | – | |
| 10 | 8.23 | – | |
| 20 | 13.58 | – | |
| 40 | – | 11.99 | |
| 50 | 24.10 | – | |
| H2S | baseline | – | |
| 5 | 0.17 | – | |
| 10 | 2.42 | – | |
| 20 | 3.87 | – | |
| 50 | 24.44 | – | |
| 80 | – |
Table 8 reports the LRT detection significance for each amplification scenario. SO2 achieves robust detection () at 5 amplification with SNR 10 and at 9 amplification () with SNR 5. NH3 requires 10 amplification at SNR 10, while H2S requires 20. CH4 is not included in the table because it remains undetected by the relative LRT (full model versus model without CH4) at all tested amplification levels. This non-detection arises from strong spectral degeneracy with H2O, whose retrieved abundance (log VMR ) produces broad near-infrared absorption that overlaps and masks the CH4 features; the retriever can compensate for the removal of CH4 by adjusting H2O, yielding a low LRT statistic even when the CH4 posterior itself is constrained (Figure 9). However, an alternative LRT formulation using a featureless (flat-line) null hypothesis confirms that the CH4 spectral signature is intrinsically detectable by BSSL, exceeding at 5 amplification, consistent with the spectral amplification results shown in Figure 8. SiO, whose spectral contribution is negligible in all scenarios, is likewise omitted from the table.
@@ 4.1 H2/He dominated exoplanets
In summary, Twinkle’s capability to characterise exoplanet atmospheres depends primarily on stellar brightness in optical and infrared wavelengths, planetary and stellar radii, main atmospheric component and cloud coverage. While stellar brightness and planetary dimensions are usually well-characterised by previous observations, cloud properties and, in the case of small planets, atmospheric mean molecular weight may require dedicated characterisation efforts. Cloud contributions may significantly impact parameters’ retrieval, especially for minor species. For instance, in our simulated retrievals of WASP-107 b, SO2 spectral features were fully obscured by the assumed cloud opacity (see Figure 5), and degeneracies between cloud-top height and planetary radius were observed. Incorporating detailed cloud microphysical and radiative models into future simulations might allow more realistic analyses.
@@ 4.3 Stellar Activity Effects, Targets Observability and Future Works
While stacking multiple observations to achieve higher SNR per wavelength bin is an effective strategy to enhance target observability, variability between observations introduces potential challenges. Variations in planetary atmospheric conditions between different transits or eclipses can lead to differences in the collected spectra. More significantly, stellar activity may significantly contaminate atmospheric parameters such as molecular abundances and temperature retrievals. A population study of 20 exoplanets conducted with HST by Saba et al. (2025) demonstrated that stellar contamination could alter molecular abundance estimates by up to six orders of magnitude and atmospheric temperatures by up to 145%. Stellar contamination predominantly affects the optical wavelengths, which are partially covered by Twinkle. Hence, the potential influence of stellar activity must be carefully considered, especially when multiple observations are stacked. Investigating stellar activity effects on exoplanet atmospheres is one of Twinkle’s core scientific themes, and a future study will explore Twinkle’s capability to identify and quantify these stellar variations. In addition to astrophysical variability, instrumental systematics such as detector drift, thermal variations, and pointing jitter may affect the consistency of stacked observations. A detailed assessment of these instrumental effects will be presented in a companion study focusing on time-domain simulations.
@@ 5 Conclusions
Our analysis indicates that approximately 100 exoplanets are suitable for in depth atmospheric studies and over 300 exoplanets are good candidates for characterisation with Twinkle.
@@ Acknowledgements
We gratefully acknowledge insightful discussions and inputs from members of the Twinkle Science Team, and their participation in Twinkle Science Conference. This work made use of data and results from observations with the NASA/ESA/CSA James Webb Space Telescope, as reported in the published literature. The JWST data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST.
@@
Conflict of Interest
The authors declare no conflict of interest.
Appendix A Noise Contributions
The noise component plots presented in this section illustrate how the different noise sources considered by the Radiometric Tool vary across each target. For all targets simulated in this work, the noise budget remains in the photon-noise-dominated regime, with all noise contributions increasing towards the edges of each spectral channel. For relatively bright targets such as HD 209458 b and 55 Cnc e, read noise begins to exceed photon noise at the longer wavelengths covered by channel 1, as the detector requires more reads to avoid saturation.
The radiometric simulations presented in this study adopt the instrument parameters summarised in Table 9. In its current design, the telescope has an effective collecting area of m2. The optical train consists of four mirror surfaces and a dichroic beam splitter in the common path, followed by a dispersing prism and three fold mirrors in each channel. All mirror surfaces are assumed to have a reflectivity of 98%, while the dichroic and prism each transmit 90% and 80%, respectively. The detector quantum efficiency is wavelength-dependent, ranging from 75% to 95% across the 0.5–4.5 m bandpass.
The point spread function (PSF) in channel 0 is treated as diffraction-limited, while channel 1 assumes a wavefront error of m. Pixel-based PSF models are used for both channels. The detectors operate at 90 K with a read noise of 25 counts per read, a dark current of 35 counts s-1, and a well depth of 72,000 counts (80% of which is usable to minimise non-linear detector effects). For bright targets such as HD 209458 b and 55 Cnc e, shorter integration times are required to avoid well saturation, leading to a higher number of detector reads per exposure and consequently elevated read noise contributions (Figures 10(a) and 11(b)). For fainter targets (WASP-107 b, GJ 3470 b), longer integrations are feasible, and photon noise dominates across the full spectral range.
Line-of-sight pointing jitter is included in the noise model using a pre-computed jitter profile derived from ground-test pointing performance data. The out-of-transit noise overhead is accounted for by a factor , corresponding to an out-of-transit baseline equal to 1.5 times the transit duration. The zodiacal background is modelled using a position-dependent sky map with a nominal scaling factor of 2.5. Additional systematic noise contributions (e.g. gain noise, detector drift, thermal variations) are not included in the present radiometric model but will be addressed in a forthcoming time-domain simulation study.
| Parameter | Value |
| Collecting area () | 0.146 m2 |
| Channel 0 range | 0.5–2.43 m () |
| Channel 1 range | 2.43–4.5 m () |
| Mirror reflectivity (per surface) | 98% |
| Dichroic transmission | 90% |
| Prism transmission | 80% |
| Detector QE | 75–95% |
| Read noise | 25 counts |
| Dark current | 35 counts s-1 |
| Well depth | 72,000 counts |
| Detector temperature | 90 K |
| Optics temperature | 100–160 K |
| PSF (Ch 0) | Diffraction-limited |
| PSF (Ch 1) | m |
The observation stacking equation (Equation 1) assumes that the noise contributions from successive transits or eclipses are uncorrelated. While this assumption is appropriate for photon noise and random detector noise, systematic effects such as instrumental drift, thermal cycling, or stellar variability may introduce correlated noise components that do not average down as . Experience with HST/WFC3 transit spectroscopy suggests that systematic noise floors of 20–30 ppm can be reached for bright targets after careful detrending (Deming et al., 2013). A detailed assessment of Twinkle’s systematic noise floor, incorporating time-domain instrumental effects, will be presented in a companion study.
Appendix B Retrieval Posterior Plots
@@
Appendix C Mass Estimates for TESS Object of Interests
In order to estimate masses for TOIs with radius available, we utilised Forecaster tool developed by Chen and Kipping (2017), a code that predict mass from radius for objects. We take the maximum radius of known exoplanets as an edge, and treat those with larger radii as dwarfs and ignore them in this study. Mass estimate using Forecaster tool for TOIs with radius less than 8 Earth radii is being adopted as they align with the known exoplanets distribution as shown in Figure 16. For TOIs with higher radius, we assume an asymmetric log-normal distribution to randomly set the mass, with the mean value set as the median value of known exoplanets with radius larger than 8 Earth radii. Resultant mass distribution for TOIs in Twinkle’s FoR is shown in Figure 16.
Appendix D Likelihood Ratio Test for Molecular Detectability
To quantify the statistical significance of molecular detections, we employ the Likelihood Ratio Test (LRT) following the methodology described by Tessenyi et al. (2013). For each molecule under consideration, the alternative hypothesis corresponds to the full forward model containing all retrieved species, while the null hypothesis corresponds to the model with the test molecule removed. For each hypothesis, Monte Carlo simulations with 100,000 iterations generate synthetic observations by adding Gaussian noise (drawn from the per-bin radiometric uncertainties) to the corresponding model spectrum. For each realisation, the likelihood ratio statistic is computed, yielding distributions of (data generated under ) and (data generated under ). A molecule is classified as a “Detection” when median() mean() , where is the standard deviation of the null distribution. The detection significance, expressed in units of , is reported in the results tables. Molecules falling below the threshold are classified as “Upper Limit”, indicating that their abundances are either prior-dominated or only weakly constrained by the data.
For emission spectra (55 Cnc e), we adopt the same molecule-removal framework: is the full emission model containing all retrieved species, while is the full model with the test molecule removed. This approach is more conservative than the original Tessenyi et al. (2013) emission formulation (which uses a featureless blackbody as ), because it tests whether a molecule’s spectral signature is distinguishable in the presence of all other atmospheric constituents, rather than merely deviating from a smooth continuum.
To ensure robustness against unreliable MAP estimates for poorly constrained parameters, we perform the LRT using both the MAP and posterior median values for each molecule. A molecule is classified as a “Detection” only if both tests independently exceed the threshold. The more conservative (lower) of the two significance values is reported in the results tables.
Figures 17–20 show the resulting (black) and (green) distributions for each molecule tested. The red dashed line indicates the detection threshold, while the blue solid line marks the median of . A molecule is detected when the blue line exceeds the red line.